OpenStudy (anonymous):

what shape is 4x^2 - 16^y = 16

OpenStudy (abb0t):

parbola.

OpenStudy (anonymous):

what if it was 16x^2 + 4y^2 = 16 instead? Would it be the same or are they trying to trip me up?

OpenStudy (anonymous):

I think this would be an ellipse, right?

OpenStudy (anonymous):

for both? or only the second one?

OpenStudy (abb0t):

wait, what??

OpenStudy (abb0t):

\(4x^2+16^y=16\) is that it?>

OpenStudy (anonymous):

There are two: 4x^2 - 16y^2 = 16 16x^2 + 4y^2 = 16 Would they be the same shape or not, due to the negative?

OpenStudy (dumbcow):

haha no that cant be it its the diff between an ellipse (+) and a hyperbola(-)

OpenStudy (abb0t):

yeah, hyperbola is the -

OpenStudy (anonymous):

Ok thank you!

OpenStudy (dumbcow):

just for fun http://www.wolframalpha.com/input/?i=plot+4x%5E2+-16%5Ey+%3D16